A pair of nuclei for which Z1 = N2 and Z2 = N1 are called mirror isobars (the atomic and neutron numbers are interchanged). Binding-energy measurements on these nuclei can be used to obtain evidence of the charge independence of nuclear forces (that is, proton
Answer #1

A way to compute the binding energies is from atomic mases of the isotopes (1 amu =931.5 MeV)

(In case of Li one can not apply the semiempircal relation for nuclear binding energy, becasue it has an anomaly)

M(7,4)Be =7.0169 amu =6536.2424 MeV (http://en.wikipedia.org/wiki/Isotopes_of_beryllium)

M(7,3)Li =7.0160 amu =6535.404 MeV    (http://en.wikipedia.org/wiki/Isotopes_of_lithium)

M(1,0)n = 1.00866 amu =939.565 MeV for neutron (see wiki)

M(1,1)p = 1.00728 amu =938.272 MeV for proton (see wiki)

The binding energies are

BE(Be) = 4*938.272 +3*939.565 - 6536.2424 = 35.54 MeV

BE(Li) = 3*938.272 +4*939.565 -6535.404 = 37.672 MeV

The difference in binding energies is

Delta(BE) =37.672-35.54 = 2.132 MeV
Answer #2

You cannot calculate the difference in binding energy with just the information above.

Now there are many ways to get the answer.

You can:-

- Check out a nuclear data table where the binding energies are listed. You would calculate the difference (simply subtract the small one from the bigger one) then divide the result by their mass number which is 7 in your question.

- you can use the masses from a nuclear chart to calculate the binding energies of each the also as above calculate the difference and divide by the mass number (which is 7)

- you can use a nuclear model and calculate the binding energies from primary principles.

- or you can calculate the difference in binding energy from the decay energy and the mass of an electron a neutron and a proton.

Hope I Helped! :-)